lata’s age after 7 years will be 40.what was her age after y years ago?

Grade:12th pass

3 Answers

Arun
5 years ago
Dear student
 
Let present age of Lata = x
 
Hence
 
x + 7 = 40
x = 33 years
.so present age = 33 years.
 
Now age y years ago = 33-y
Crazy learner
5 years ago
Let Laura's present are be x
Lata's age after 7 years is 40,which gives us the following equation:
                                                    x+7=40
                                                     x=33 (lata's present age)
Her age y years ago will be 33-y
Khimraj
5 years ago
Let persent age of  lata is x
then
X + 7 = 40
x = 33
and age before y year ago = 33 -y.
Hopr it clears.